# Advances on the finite element discretization of fluid-structure interaction problems

### v1

@@ 2.1. The continuous problem
$\mathbf{X}(\mathbf{s},0)=\mathbf{X}_{0}(\mathbf{s})$ $\quad\forall\mathbf{s}\in\mathcal{B},$
where the following bilinear and trilinear forms are used in the fluid domain
$a(\mathbf{u},\mathbf{v})=\left(\nu\operatorname*{{\underline{\bm{\varepsilon}}}}(\mathbf{u}),\operatorname*{{\underline{\bm{\varepsilon}}}}(\mathbf{v})\right)_{\Omega},\qquad b(\mathbf{u},\mathbf{v},\mathbf{w})=\frac{\rho_{f}}{2}\left((\mathbf{u}\cdot\nabla\mathbf{v},\mathbf{w})_{\Omega}-(\mathbf{u}\cdot\nabla\mathbf{w},\mathbf{v})_{\Omega}\right).$
The extended viscosity $\nu$ equals $\nu_{f}$ in $\Omega^{f}_{t}$ and $\nu_{s}$ in $\Omega^{s}_{t}$. Moreover, $\delta\rho=\rho_{s}-\rho_{f}$.
The reader interested in the detailed derivation of the above problem may refer to [17].
Remark 8.
Existence and uniqueness of the solution for Problem 4 were proved in [24] for a simplified version of the problem by following the Galerkin approximation technique used in [72]. In [24], we considered a linearized version of the problem: the convective term was neglected and a linear constitutive law was chosen for the solid, that is $\mathbb{P}(\mathbb{F})=\kappa\mathbb{F}$. Moreover, the coupling form $\mathbf{c}(\cdot,\cdot)$ was set to be the scalar product in $\mathbf{H}^{1}(\mathcal{B})$. We proved the following space-time regularity
@@
$\mathscr{H}_{0}=\text{ the closure of }\mathscr{V}\text{ in }\mathbf{H}^{1}_{0}(\Omega),\qquad\mathscr{V}_{0}=\text{ the closure of }\mathscr{V}\text{ in }\mathbf{L}_{0}^{2}(\Omega),$
@@ 2.2. Discretization
We discretize Problem 4 by mixed finite elements.
Similarly to what we did for the elliptic interface problem, fluid and solid (reference) domain are treated independently. We thus partition $\Omega$ by a mesh $\mathcal{T}_{h}^{\Omega}$ with size $h_{\Omega}$, while we decompose the domain $\mathcal{B}$ by a mesh $\mathcal{T}_{h}^{\mathcal{B}}$ with size $h_{\mathcal{B}}$. We emphasize that the considered meshes are fixed throughout the entire evolution of the system.
We then choose four finite dimensional subspaces, ${\mathbf{V}_{h}\subset\mathbf{H}^{1}_{0}(\Omega)}$, ${Q_{h}\subset L^{2}_{0}(\Omega)}$, ${\mathbf{S}_{h}\subset\mathbf{H}^{1}(\mathcal{B})}$ and ${{\bm{\Lambda}}_{h}\subset{\bm{\Lambda}}}$. In order to design a stable method, the velocity and pressure spaces $(\mathbf{V}_{h},Q_{h})$ must be a compatible pair for the Stokes problem [15]. Compatibility requirements involving the solid discrete spaces $\mathbf{S}_{h}$ and ${\bm{\Lambda}}_{h}$ have been previously established for the elliptic interface problems and apply straightforwardly to this case. For more details, see Section 1.2 and the discrete spaces therein.
@@
$\mathbf{u}_{s,0}=\frac{\mathbf{X}_{h}^{0}-\mathbf{X}_{h}^{-1}}{\Delta t}\qquad\text{in }\mathcal{B}.$
@@
In any case, one of the important features of our semi-implicit scheme is that it is unconditionally stable, as stated by the following proposition.
Proposition 5 ([17]).
The discrete Problem 5 is unconditionally stable in time, without any restriction on the choice of time step $\Delta t$.
Remark 9.
Higher order schemes in time may also be considered in the spirit of [43, 65, 36]. In [26], we proved that the second order backward differentiation formula BDF2 gives an unconditionally stable scheme. The same result is true also for the Crank–Nicolson scheme implemented with the midpoint rule. A proof of unconditional stability is not available for the trapezoidal Crank–Nicolson, but several numerical tests did not show any kind of instability.
We rewrite Problem 5 in matrix form by considering the special case of a linear constitutive law for the solid, that is $\mathbb{P}(\mathbb{F})=\kappa\mathbb{F}$. In the following, $\bm{\phi}$, $\psi$, $\bm{\chi}$, $\bm{\zeta}$ denotes the basis functions for $\mathbf{V}_{h}$, $Q_{h}$, $\mathbf{S}_{h}$ and ${\bm{\Lambda}}_{h}$, respectively. We have
@@
The coupling matrix $\mathsf{C}_{f}(\mathbf{X}_{h}^{n})$ corresponds to the term $\mathbf{c}(\bm{\mu}_{h},\mathbf{v}_{h}(\mathbf{X}_{h}^{n}))$ and represents the interaction between the fluid and the immersed solid. We point out that its construction requires the integration over $\mathcal{B}$ of $\mathbf{v}_{h}$ (defined on $\mathcal{T}_{h}^{\Omega}$) composed with the deformation map $\mathbf{X}_{h}^{n}$: this takes care of the actual position of the solid body with respect to the background mesh, as shown in Figure 7. It is clear that the assembly of $\mathsf{C}_{f}(\mathbf{X}_{h}^{n})$ falls in the framework we described in Section 1.3.
In the case of a nonlinear solid constitutive law, the matrix $\mathsf{A}_{s}$ depends on the deformation $\mathbf{X}$, thus the system must be solved by employing a solver for nonlinear systems of equations, such as a fixed point iteration or a Newton-like method.
At each time step, the system (27) can be interpreted as a stationary saddle point problem, which admits a unique solution. Stability and well-posedness of its finite element discretization were proved in [23]: the method gives optimal convergence rates, which depend on the regularity of the solution. Well-posedness and convergence properties in case of inexact integration of the coupling term were discussed in [19].
@@ 2.3.1. Convergence analysis for stationary problem
We consider a stationary version of Problem 5 and we study how the convergence of the method is affected by the assembly technique for the coupling term. Here we discuss the example of the immersed square presented in [19]. The interested reader can refer to [18, 19] for a wider set of numerical tests.
@@
(28) $\mathbf{u}(x,y)=\operatorname{curl}\left((4-x^{2})^{2}(4-y^{2})^{2}\right),$ $\quad p(x,y)=150\,\sin(x),$
@@ 2.3.2. The stretched annulus
[…]
@@
The equations are discretized by using the $(\mathcal{Q}_{2},\mathcal{P}_{1},\mathcal{Q}_{1},\mathcal{Q}_{1})$ finite element. The linear system is solved using GMRES accelerated by the preconditioners introduced in Section 1.4, whose action is performed through exact inversion. The performance of the parallel solver has been analyzed, in terms of optimality and scalability, in [22]. The coupling term is the scalar product in $L^{2}(\mathcal{B})$ assembled exactly.
@@ 3. Conclusions
Due to jumps in the coefficient or the immersion of the solid body in the fluid, the physical domain is partitioned into two regions. The main idea behind our approach is to extend one domain into the other. Then, the equations are solved on both the background and the immersed domains, while a coupling term describes their interaction in the fictitiously extended region. The computational domains are fixed and discretized by two independent meshes.
After recalling the main stability properties of our approach, we discussed how to deal with the coupling term. The computation of its discrete version involves integration of functions defined over two distinct non-matching grids. The coupling term can be computed exactly by considering a composite quadrature rule on the intersection between the two meshes, or in a (cheaper) approximate way taking into account the presence of an additional error.

### v2

@@ 2.1. The continuous problem
$\mathbf{X}(\mathbf{s},0)=\mathbf{X}_{0}(\mathbf{s})$ $\quad\forall\mathbf{s}\in\mathcal{B}.$
The reader interested in the detailed derivation of the above problem may refer to [37]. We define the following bilinear and trilinear forms on the fluid domain
$a(\mathbf{u},\mathbf{v})=\left(\nu\operatorname*{{\underline{\bm{\varepsilon}}}}(\mathbf{u}),\operatorname*{{\underline{\bm{\varepsilon}}}}(\mathbf{v})\right)_{\Omega},\qquad b(\mathbf{u},\mathbf{v},\mathbf{w})=\frac{\rho_{f}}{2}\left((\mathbf{u}\cdot\nabla\mathbf{v},\mathbf{w})_{\Omega}-(\mathbf{u}\cdot\nabla\mathbf{w},\mathbf{v})_{\Omega}\right),$
where the extended viscosity $\nu$ equals $\nu_{f}$ in $\Omega^{f}_{t}$ and $\nu_{s}$ in $\Omega^{s}_{t}$. Moreover, $\delta\rho=\rho_{s}-\rho_{f}$.
Remark 2.2.
Existence and uniqueness of the solution for Problem 2.1 have been proved in [78] for a simplified version of the problem by following the Galerkin approximation technique used in [79]. In [78], a linearized version of the problem is considered: indeed the convective term is neglected and a linear constitutive law is chosen for the solid, that is $\mathbb{P}(\mathbb{F})=\kappa\mathbb{F}$. Moreover, the coupling form $\mathbf{c}(\cdot,\cdot)$ is set to the scalar product in $\mathbf{H}^{1}(\mathcal{B})$. The following space-time regularity is ensured
@@
[…]
@@ 2.2. Discretization
We discretize Problem 2.1 by mixed finite elements.
Similarly to what we did for the elliptic interface problem, fluid and solid (reference) domain are treated independently. We thus partition $\Omega$ by a mesh $\mathcal{T}_{h}^{\Omega}$ with size $h_{\Omega}$, while we decompose the domain $\mathcal{B}$ by a mesh $\mathcal{T}_{h}^{\mathcal{B}}$ with size $h_{\mathcal{B}}$. We emphasize that the considered meshes are fix throughout the entire evolution of the system.
We then choose four finite dimensional subspaces, namely $\mathbf{V}_{h}\subset\mathbf{H}^{1}_{0}(\Omega)$}, $Q_{h}\subset L^{2}_{0}(\Omega)$}, $\mathbf{S}_{h}\subset\mathbf{H}^{1}(\mathcal{B})$} and ${\bm{\Lambda}}_{h}\subset{\bm{\Lambda}}$}. In order to design a stable method, the velocity and pressure spaces $(\mathbf{V}_{h},Q_{h})$ must be a compatible pair for the Stokes problem [59]. Compatibility requirements involving the solid discrete spaces $\mathbf{S}_{h}$ and ${\bm{\Lambda}}_{h}$ have been previously established for the elliptic interface problems and apply straightforwardly to this case. For more details, see Section 1.2 and discrete spaces therein.
@@
[…]
@@
In any case, our semi-implicit scheme is unconditionally stable, as stated by the following proposition.
Proposition 2.4 ([37]).
The discrete Problem 2.3 is unconditionally stable in time, without any restriction on the choice of time step $\Delta t$.
Remark 2.5.
We chose Backward Euler for the time discretization, but high order schemes may also be considered in the spirit of [80, 81, 82]. In [83], we proved that the second order backward differentiation formula BDF2 gives an unconditionally stable scheme. The same result is also true for the Crank–Nicolson scheme implemented with the midpoint rule. A proof of unconditional stability is not available for the trapezoidal Crank–Nicolson, but several numerical tests did not show any kind of instability.
Remark 2.6.
The incompressibility of both fluid and solid is handled by imposing the divergence-free constraint on the extended velocity $\mathbf{u}$, and guarantees conservation of mass at the continuous level. Conversely, at the discrete level, the divergence-free constraint may not be exactly imposed. Indeed, it is well-known that many popular mixed finite element pairs for the Stokes equation (Hood–Taylor, Bercovier–Pironneau, MINI, etc., see [59]) conserve mass only in approximate way [84], while the construction of conforming divergence-free finite elements is not trivial and requires special recipes (see [85, 86, 87], for instance). The mass conservation properties of our approach in terms of conforming finite elements have been discussed in [88, 89] and further details will be addressed in forthcoming works.
We rewrite Problem 2.3 in matrix form by considering the special case of a linear constitutive law for the solid, that is $\mathbb{P}(\mathbb{F})=\kappa\mathbb{F}$. In the following, $\bm{\phi}$, $\psi$, $\bm{\chi}$, $\bm{\zeta}$ denotes the basis functions for $\mathbf{V}_{h}$, $Q_{h}$, $\mathbf{S}_{h}$ and ${\bm{\Lambda}}_{h}$, respectively. We have
@@
The coupling matrix $\mathsf{C}_{f}(\mathbf{X}_{h}^{n})$ corresponds to the terms of kind $\mathbf{c}(\bm{\mu}_{h},\mathbf{v}_{h}(\mathbf{X}_{h}^{n}))$ and represents the interaction between the fluid and the immersed solid. We point out that its construction requires the integration over $\mathcal{B}$ of $\mathbf{v}_{h}$ (defined on $\mathcal{T}_{h}^{\Omega}$) composed with the deformation map $\mathbf{X}_{h}^{n}$: this takes care of the actual position of the solid body with respect to the background mesh, as sketched in Figure 7. It is clear that the assembly of $\mathsf{C}_{f}(\mathbf{X}_{h}^{n})$ falls in the framework we described in Section 1.3.
In the case of a nonlinear solid constitutive law, the matrix $\mathsf{A}_{s}$ depends on the deformation $\mathbf{X}$, thus the system must be solved by employing a solver for nonlinear systems of equations, such a fixed point iteration or a Newton-like method.
At each time step, system (26) can be interpreted as a stationary saddle point problem, which admits a unique solution. Stability and well-posedness of the finite element discretization have been proved in [69]: the method gives optimal convergence rates, which depend on the regularity of the solution. Well-posedness and convergence properties in case of inexact integration of the coupling term have been discussed in [39].
@@ 2.3.1. Convergence analysis for stationary problem
We consider a stationary version of Problem 2.3 and we study how the convergence of the method is affected by the assembly technique for the coupling term. Here we only discuss the example of the immersed square presented in [39]. The interested reader can refer to [38, 39] for a wider set of numerical tests.
@@
[…]
@@ 2.3.2. The stretched annulus
The solid material is modeled by the linear constitutive law $\mathbb{P}(\mathbb{F})=\kappa\mathbb{F}$, with $\kappa=10$. At rest, the annulus occupies the region $\{\mathbf{x}\in\mathbb{R}^{2}:\,0.3\leq|\mathbf{x}|\leq 0.5\}$, being at the center of the box $1,1]^{2}$ filled by the fluid. Since the geometry of the system is symmetric, we can reduce our study to a quarter of the domain, thus we set
@@
The equations are discretized by the $(\mathcal{Q}_{2},\mathcal{P}_{1},\mathcal{Q}_{1},\mathcal{Q}_{1})$ finite element. The linear system is solved using GMRES accelerated by the preconditioners introduced in Section 1.4, whose action is performed through exact inversion. The performance of the parallel solver has been analyzed, in terms of optimality and scalability, in [45]. The coupling term is the scalar product in $L^{2}(\mathcal{B})$ assembled exactly.
@@ 3. Conclusions
Due to jumps in the coefficient or the immersion of the solid body in the fluid, the physical domain is partitioned into two regions. The main idea behind our approach is to extend one domain to the other. Then, the equations are solved on both the background and the immersed domains, while a coupling term describes their interaction in the fictitiously extended region. The computational domains are fixed and discretized by two independent meshes.
After recalling the main stability properties of our approach, we discussed how to deal with the coupling term. Indeed, the computation of its discrete version involves integration of functions defined over two distinct non-matching grids. The coupling term can be computed exactly by considering a composite quadrature rule on the intersection between the two meshes, or in a (cheaper) approximate way taking into account the presence of an additional error.